Problem 9-4A

At the beginning of 2013, Mazzaro Company acquired equipment costing $127,600. It was estimated that this equipment would have a useful life of 6 years and a salvage value of $12,760 at that time. The straight-line method of depreciation was considered the most appropriate to use with this type of equipment. Depreciation is to be recorded at the end of each year.

During 2015 (the third year of the equipment’s life), the company’s engineers reconsidered their expectations and estimated that the equipment’s useful life would probably be 7 years (in total) instead of 6 years. The estimated salvage value was not changed at that time. However, during 2018, the estimated salvage value was reduced to $5,317.

Indicate how much depreciation expense should be recorded each year for this equipment, by completing the following table. (Round answers to 0 decimal places, e.g. 5,275.)


Year

Depreciation
Expense

Accumulated
Depreciation

2013 $

$

2014

2015

2016

2017

2018

2019
Answer #1

Indicate how much depreciation expense should be recorded each year for this equipment, by completing the following table. (Round answers to 0 decimal places, e.g. 5,275.)


Year

Depreciation
Expense

Accumulated
Depreciation

2013 (127600-12760/6) = 19140

$19140

2014

19140

38280

2015

(127600-38280-12760/5) = 15312

53592

2016

15312

68904

2017

15312

84216

2018

(127600-84216-5317)/2 = 19034

103250

2019

19033

122283